QUESTION 7 A centrifugal pump discharges 500 GPM through a 12-inch diameter impeller. The fluid leaves the impeller with a taPlease answer question 7:

A centrifugal pump discharges 500 GPM through a 12-inch diameter impeller.  The fluid leaves the impeller with a tangential velocity of 17 ft/sec.  If water, @ 68°F, enters the impeller via radial flow (I.E. no whirl), and the inner radius and thickness of the impeller are 1.75-inches.  (Assume a constant thickness of the impeller) and if the pump speed is 1775 rpm, with a head loss of 4.5-feet.  What is the horsepower input to the flow by the pump?
Answer #1

N . Given, Discharge Q = 500 GPM 1 = 1.114 t/sec Diameter of impeller d = gets 12 = 1 Diesen Tangential velocity (42) = 17= na Mm = 162.43) (1114) (9.8856) = 687-514 lb . int-sec = 687.514 550 = 1:25002 ~ 1.25HP (b)